Using the Straight Line (SL) depreciation method, what is the (a) annual depreciation charge, Dt=?, (b) the total accumulated depreciation charge, D3=?, at the end of year 3, and (c) the Book Value, BV3=?, at the end of year 3, for an automobile with a five-year useful life that has a first cost of $20,200 and an $8,565.00 salvage value?
Answer #1

Depreciation per year = (P-S)/N = (20200 - 8565) / 5 = 2327

a. Depreciation in year 3 = 2327

b. Accumulated depreciation in year 3 = 2327 * 3 = 6981

c. BV at end of yr 3 = 20200 - 6981 = 13219